You are installing a solar panel today on the roof of your house and it costs 4000. Over the next 12 months the expected savings from this investment can be either 300 or 100 with equal probability. For simplicity you can assume that those savings will be realized in perpetuity. The opportunity cost of capital is 5%. Is this a good investment today? What about if you wait for a year?
Answer #1

Cost of solar panel =$4000
Weighted average cost of saving = (300+100)/2 = 200

The present value of cash flow from perpetuity= cost of saving÷cost of capital
= 200÷5%
=$4000

Is this a good investment today?
It is neither a good or bad investment. As the present value of savings= cost of the solar panel.

Even after one year, the answer remains the same, as cash flow is till perpetuity. [Present value of cash flow from perpetuity= cost of saving÷cost of capital
= 200÷5%
=$4000]